# JaxMARL: Multi-Agent RL Environments and Algorithms in JAX

### v4

@@ Abstract.
Benchmarks play an important role in the development of machine learning algorithms. For example, research in reinforcement learning (RL) has been heavily influenced by available environments and benchmarks. However, RL environments are traditionally run on the CPU, limiting their scalability with typical academic compute. Recent advancements in JAX have enabled the wider use of hardware acceleration to overcome these computational hurdles, enabling massively parallel RL training pipelines and environments. This is particularly useful for multi-agent reinforcement learning (MARL) research. First of all, multiple agents must be considered at each environment step, adding computational burden, and secondly, the sample complexity is increased due to non-stationarity, decentralised partial observability, or other MARL challenges. In this paper, we present JaxMARL, the first open-source code base that combines ease-of-use with GPU enabled efficiency, and supports a large number of commonly used MARL environments as well as popular baseline algorithms. When considering wall clock time, our experiments show that per-run our JAX-based training pipeline is up to 12500x faster than existing approaches. This enables efficient and thorough evaluations, with the potential to alleviate the evaluation crisis of the field. We also introduce and benchmark SMAX, a vectorised, simplified version of the popular StarCraft Multi-Agent Challenge, which removes the need to run the StarCraft II game engine. This not only enables GPU acceleration, but also provides a more flexible MARL environment, unlocking the potential for self-play, meta-learning, and other future applications in MARL. We provide code at https://github.com/flairox/jaxmarl.
Key words and phrases:
Multi-Agent Reinforcement Learning, JAX, Benchmarks1. Introduction
Benchmarks play a pivotal role in the development of new single and multi-agent reinforcement learning (MARL) algorithms by defining problems, enabling comparisons, and focusing efforts. For example, in recent years, Go and Chess drove the development of MuZero Schrittwieser et al. (2020) while decentralised StarCraft Micromanagement Foerster et al. (2017) and later the StarCraft Multi-Agent Challenge (SMAC, Samvelyan et al., 2019) resulted in the development of algorithms such as QMIX Rashid et al. (2020), a popular MARL technique.
Data transfer between the CPU (where the environment is simulated) and the GPU (where the agents are evaluated) is a crucial bottleneck for simulation speed. Simulation speed in turn is vital for progress in reinforcement learning (RL) because RL algorithms often require a large number of environment interactions. This problem is even worse in MARL, where non-stationarity and decentralised partial observability greatly worsen the sample complexity Bernstein et al. (2002). Hardware acceleration and parallelisation are crucial to alleviating this, but current acceleration and parallelisation methods are typically not implemented in Python, reducing their accessibility for most machine learning researchers Shacklett et al. (2023); Weng et al. (2022). For example, the extremely efficient Hanabi library Hu and Foerster (2020) from Meta-AI research is implemented in C++ and has seen relatively little adoption by the community. However, recent advances in JAX Bradbury et al. (2018) have opened up new possibilities for using Python code directly with hardware accelerators, enabling the wider use of massively parallel RL training pipelines and environments.
The JAX Bradbury et al. (2018) library provides composable function transformations, allowing for automatic vectorisation, device parallelisation, automatic differentiation and just-in-time (JIT) compilation with XLA Sabne (2020), for device-agnostic optimisation. Using JAX, both the environment rollouts and model training can happen on a hardware accelerator (such as a GPU or TPU), removing the cost of data transfer between devices and allowing for significant parallelisation. Recently, PureJaxRL Lu et al. (2022a, 2023b) has demonstrated the power of this end-to-end JAX-based approach; running both the environment and the model training on a GPU yields a 4000x speedup over a “traditional” pipeline with a GPU-trained policy but a CPU-based environment.
These accelerations could substantially advance RL and MARL research by quickening the testing and iteration of ideas. Furthermore, they lower computational hurdles for in-depth MARL research, enabling researchers to utilise billions of frames and extract more performance from single GPUs.
@@
Alongside the current computational issues faced by MARL researchers, recent work also highlights issues with the evaluation standards and use of benchmarks in the MARL community. In particular, MARL papers typically only test on a few domains. Of the 75 recent MARL papers analysed by Gorsane et al. (2022), 50% used only one evaluation environment and a further 30% used only two. While SMAC and MPE Lowe et al. (2017), the two most used environments, have various tasks or maps, the lack of a standard set raises the risk of biased comparisons and incorrect conclusions. This leads to environment overfitting and unclear progress markers.
Instead, novel MARL methods should be tested on a wide range of domains to accurately evaluate their limits and enable better comparisons. The likely issue preventing this is the lack of a unified codebase and the computational burden of further evaluation.
This paper presents JaxMARL, a Python library that for the first time brings together JAX implementations of eight common MARL environments under one API. We additionally provide JAX implementations for four state-of-the-art algorithms, allowing for end-to-end JAX-based training pipelines in a similar fashion to PureJaxRL. As outlined in Figure 1, we present a library with end-to-end hardware-accelerated training, simple Python implementations, and a broad range of MARL environments. By alleviating computational constraints, JaxMARL allows rapid evaluation of novel methods across a broad set of domains, and hence has the potential to be a powerful tool to address MARL’s evaluation crisis. Specifically, we find that JaxMARL achieves over 12500x speedup compared to “conventional” aproaches.
We also create SMAX, a JAX-based simplification of the centralised training with decentralised execution (CTDE) benchmarks SMAC Samvelyan et al. (2019) and SMACv2 Ellis et al. (2022). SMAX features simplified dynamics, greater flexibility and a more sophisticated but fully-decentralised heuristic AI, while retaining the high-dimensional observation space, complex unit type interactions and procedural scenario generation that lend SMAC and SMACv2 much of their difficulty.
As shown in Figure 2, in addition to SMAX, our library includes the most popular environments from several MARL settings. For centralised training with decentralised execution (CTDE), we include the Multi-Agent Particle Environments (MPE) Lowe et al. (2017), and Multi-Agent Brax (MABrax). Meanwhile, for zero-shot coordination (ZSC) and ad-hoc teamplay, we include Hanabi and Overcooked. Lastly, from the general-sum literature, we include the CoinGame and Spatial-Temporal Representations of Matrix Games (STORM), a representation of matrix games as grid-world scenarios with temporally extended actions. JaxMARL provides the first JAX implementation of these environments and unifies them in a single codebase.
We additionally provide JAX implementations of Independent PPO (IPPO) Schulman et al. (2017); de Witt et al. (2020), QMIX, VDN Sunehag et al. (2017) and Independent -Learning (IQL) Mnih et al. (2015), four of the most common MARL algorithms, allowing new techniques to be easily benchmarked against existing practices. We will extend this list before the camera-ready copy, e.g. with the popular MAPPO Yu et al. (2022) algorithm.
2. Background
2.1. Hardware Accelerated Environments
@@
JAX enables the use of Python code with any hardware accelerator, allowing researchers to write hardware-accelerated code easily. Within the RL community, writing environment code in JAX has gained recent popularity. This brings two chief advantages: firstly, environments written in JAX can be very easily parallelised by using JAX’s vmap operation, which vectorises a function across an input dimension, and secondly writing the environment in JAX allows the agent and environment to be co-located on the GPU, which eliminates the time taken to copy between CPU and GPU memory. Combined, these two factors bring significant increases in training speed, with PureJaxRL Lu et al. (2022a) achieving a x speedup over traditional training in single-agent settings.
2.2. SMAC
StarCraft is a popular environment for testing RL algorithms. It typically features features a centralised controller issuing commands to balance micromanagement, the low-level control of individual units, and macromanagement, the high level plans for economy and resource management.
SMAC Samvelyan et al. (2019), instead, focuses on decentralised unit micromanagement across a range of scenarios divided into three broad categories: symmetric, where each side has the same units, asymmetric, where the enemy team has more units, and micro-trick, which are scenarios designed specifically to feature a particular StarCraft micromanagement strategy. SMACv2 Ellis et al. (2022) demonstrates that open-loop policies can be effective on SMAC and adds additional randomly generated scenarios to rectify SMAC’s lack of stochasticity. However, both of these environments rely on running the full game of StarCraft II, which severely increases their CPU and memory requirements. SMAClite Michalski et al. (2023) attempts to alleviate this computational burden by recreating the SMAC environment primarily in NumPy, with some core components written in C++. While this is much more lightweight than SMAC, it cannot be run on a GPU and therefore cannot be parallelised effectively with typical academic hardware, which commonly has very few CPU cores compared to industry clusters.
3. JaxMARL
@@
We present JaxMARL, a library containing simple and accessible JAX implementations of popular MARL environments and algorithms. JAX enables significant acceleration and parallelisation over existing implementations. To the best of our knowledge, JaxMARLis the first open source library that provides JAX-based implementations of a wide range of MARL environments and baselines.
3.1. API
The interface of JaxMARL is inspired by PettingZoo Terry et al. (2021) and Gymnax. We designed it to be a simple and easy-to-use interface for a wide-range of MARL problems. An example of instantiating an environment from JaxMARL’s registry and executing one transition is presented in Figure 3.
This key is separated from the internal state for clarity.
Thus, we do not use PettingZoo’s agent iterator.
3.2. Environments
@@
JaxMARL contains a diverse range of environments, all implemented in JAX. We also introduce SMAX, a SMAC-like environment implemented entirely in JAX. In this section we introduce these environments and provide details on their implementations.

### v5

@@ Abstract
Benchmarks are crucial in the development of machine learning algorithms, with available environments significantly influencing reinforcement learning (RL) research. Traditionally, RL environments run on the CPU, which limits their scalability with typical academic compute. However, recent advancements in JAX have enabled the wider use of hardware acceleration, enabling massively parallel RL training pipelines and environments. While this has been successfully applied to single-agent RL, it has not yet been widely adopted for multi-agent scenarios. In this paper, we present JaxMARL, the first open-source, Python-based library that combines GPU-enabled efficiency with support for a large number of commonly used MARL environments and popular baseline algorithms. Our experiments show that, in terms of wall clock time, our JAX-based training pipeline is around 14 times faster than existing approaches, and up to 12500x when multiple training runs are vectorized. This enables efficient and thorough evaluations, potentially alleviating the evaluation crisis in the field. We also introduce and benchmark SMAX, a JAX-based approximate reimplementation of the popular StarCraft Multi-Agent Challenge, which removes the need to run the StarCraft II game engine. This not only enables GPU acceleration, but also provides a more flexible MARL environment, unlocking the potential for self-play, meta-learning, and other future applications in MARL. The code is available at https://github.com/flairox/jaxmarl.
1 Introduction
Benchmarks are crucial for developing new single and multi-agent reinforcement learning (MARL) algorithms. They define problems, enable comparisons, and focus research efforts. For example, the development of MuZero was driven by the challenges presented by Go and Chess [54]. Similarly, decentralised StarCraft Micromanagement tasks [17] led to the creation of algorithms like QMIX [52], a popular MARL technique.
In RL research, the runtime of simulations and algorithms is a critical factor affecting the efficiency, thoroughness, and feasibility of experiments. RL training pipelines often require a large number of environment interactions and long, expensive, experimental runs significantly impede research progress.
Hardware acceleration and parallelization is an approach to address this: by running environments on hardware accelerators (e.g. GPUs), we can use many more environment instances in parallel than is feasible with a CPU, drastically improving runtime. However, such an approach typically requires significant engineering effort and often relies on non-Python codebases [56], which reduces its accessibility for many Machine Learning researchers where Python is the lingua franca.
That said, recent releases, such as the JAX [7] library and PyTorch’s functorch module [22], have improved accessibility by enabling Python code to be parallelized and just-in-time compiled on hardware accelerators.
This laid the foundation for PureJaxRL [38], which leveraged JAX to implement a parallelized approach, demonstrating that running both the environment and the model training on the same GPU yields a 10x speedup over a traditional pipeline with a GPU-trained policy but a CPU-based environment, and 4000x when multiple training runs are vectorized. This speedup enables new research directions [39, 28] and makes large-scale RL research more accessible [46].
We introduce JaxMARL, which brings these benefits to multi-agent learning. To the best of our knowledge, JaxMARL is the first open-source, Python-based library which leverages JAX for GPU acceleration and supports a wide range of popular MARL environments (as shown in Figure 1) as well as algorithms. We show that MARL greatly benefits from this approach as under the traditional approach, of using CPU-based environments, experiments tend to be particularly slow due to the increased computational burden of training multiple agents simultaneously and the higher sample complexity arising from challenges like non-stationarity and decentralized partial observability. Utilizing an end-to-end JAX-based pipeline for MARL significantly accelerates these experiments, opening up new possibilities for research in this field.
@@ 1 Introduction
Alongside computational issues, MARL research also struggles with thorough evaluation standards [21]. In particular, MARL papers typically only test on a few domains. Of the 75 recent MARL papers analysed by [21], 50% used only one evaluation environment and a further 30% used only two. While SMAC [53] and MPE [36], the two most used environments, have various tasks or maps, the lack of a standard set raises the risk of biased comparisons and incorrect conclusions. This leads to environment overfitting and unclear progress markers. By alleviating computational constraints, JaxMARL allows for rapid evaluations across a broad set of environments and hence is a powerful tool to address MARL’s current evaluation crisis.
More specifically, in this paper, our contributions are as follows:
JAX Implementations of Popular MARL Environments: We implement a wide range of popular MARL environments in JAX, enabling fast experimentation across diverse environments. Taking advantage of large scale parallelism, many of our environments run over three orders of magnitude faster on a GPU than their CPU-based counterparts. They are implemented in Python ensuring ease-of-use, following our philosophy set out in Figure 2.
New MARL Environment Suites: We introduce two new MARL environment suites: SMAX and STORM. SMAX is an approximate reimplementation of the popular SMAC(v2) [15] benchmark entirely in JAX, rather than using the StarCraft II game engine like SMAC. It is therefore more customizable and significantly faster: SMAX training is 40,000x faster than the equivalent SMAC implementation on a single NVIDIA 2080 when multiple training runs are vectorized. STORM is a general-sum environment suite inspired by the Melting Pot [34] environment suite that features temporally extended actions within social dilemmas.
Implementation of Popular MARL Algorithms in JAX: We implement many popular MARL algorithms in JAX, such as IPPO, MAPPO and QMIX. As outlined in Figure 3, our training pipeline is up to 14x faster than current popular approaches, and up to 12500x when multiple training runs are vectorized.
Comprehensive Benchmarking: We thoroughly benchmark the speed and correctness of our environments and algorithms, comparing them to existing popular repositories. Generally, our end-to-end JAX implementations run several thousand times faster than their CPU-based counterparts while maintaining equivalent agent performance.
Environment Evaluation Recommendations and Best Practice: Finally, we provide environment evaluation recommendations for different MARL research settings, such as centralized training with decentralized execution and zero-shot coordination. We also provide scripts for large scale evaluation and plotting based on best-practice in the field.
2 Background
Our work brings the benefits of hardware acceleration to multi-agent reinforcement learning.
Hardware Accelerated Environments
@@ 2 Background
JAX enables the use of Python code with any hardware accelerator, allowing researchers to write hardware-accelerated code easily. Within the RL community, writing environment code in JAX has gained recent popularity. This brings three chief advantages. Firstly, environments written in JAX can be very easily parallelised by using JAX’s vmap operation, which vectorises a function across an input dimension. Secondly writing the environment in JAX allows the agent and environment to be co-located on the GPU, which eliminates the time taken to copy between CPU and GPU memory. Finally, the code written can be compiled just-in-time, thereby improving performance. Combined, these factors bring significant increases in training speed, with PureJaxRL [38] achieving a x speedup with vectorized training over traditional training in single-agent settings.
Multi-Agent Reinforcement Learning Settings
Multi-Agent Reinforcement Learning is a subfield of reinforcement learning that focuses on environments where multiple agents interact and learn simultaneously. MARL encompasses various settings that address interactions between multiple agents. The most widely-studied MARL setting is the fully-cooperative one, in which agents work together to achieve a common goal. One key framework is centralized training with decentralized execution (CTDE) [37], which allows agents to share information during the learning phase or use additional environment information. During execution, however, agents act based on their independent and partial observations of the environment. Another is zero-shot coordination, which focuses on training agents to coordinate successfully with unseen partners or in new environments without additional training [30]. Beyond fully-cooperative settings, there are zero-sum and general-sum [34] benchmarks that study competitive and mixed incentive interactions.
3 JaxMARL
@@
We present JaxMARL, a library containing simple and accessible JAX implementations of popular MARL environments and algorithms. JaxMARL enables significant acceleration and parallelisation over existing implementations. To the best of our knowledge, JaxMARL is the first open-source library that provides JAX-based implementations of a wide range of both MARL environments and baselines. JaxMARL’s interface is inspired by PettingZoo [61] and Gymnax [32], while the design philosophy is based on PureJaxRL [38] and CleanRL [26]. We designed it to be a simple and easy-to-use interface for a wide range of MARL problems. A full specification is provided in the Appendix.
3.1 Environments
@@
JaxMARL contains a diverse range of JAX reimplementations of existing environments. It also introduces SMAX, a novel SMAC-like JAX environment, and STORM, an expansion of matrix games to grid-world scenarios. In this section, we introduce our environments while further details on their implementations can be found in the Appendix.
We measure the speed of our environments in steps per second when using random actions and compare our speed to that of the original environments in Table 3, see the Appendix for details.
3.1.1 New Environments
@@
The StarCraft Multi-Agent Challenge (SMAC) is a popular benchmark in cooperative multi-agent reinforcement learning (MARL) but has several limitations. SMAC’s environment lacks sufficient stochasticity for complex policies [15], and its reliance on the StarCraft II engine makes it slow and memory-intensive [43]. Additionally, StarCraft II’s constraints limit scenario variety and do not support competitive self-play without significant engineering. To address these issues, we introduce SMAX, a SMAC-like, hardware-accelerated, customizable environment. SMAX features more lightweight dynamics and a less exploitable AI. SMAX incorporates original SMAC scenarios and scenarios similar to those in SMACv2, but is also far more customizable. We provide more details on SMAX and how it improves on SMAC and SMACv2 in the Appendix.
Spatial-Temporal Representations of Matrix Games (STORM)
Inspired by the “in the Matrix” games in Melting Pot 2.0 [1], the STORM [29] environment expands on matrix games by representing them as grid-world scenarios. Agents collect resources which define their strategy during interactions and are rewarded based on a pre-specified payoff matrix. STORM can represent cooperative, competitive or general-sum games, like the prisoner’s dilemma [57]. Thus, STORM can be used for studying paradigms such as opponent shaping, where agents act with the intent to change other agents’ learning dynamics, which has been empirically shown to lead to more prosocial outcomes [18, 65, 41, 29, 69]. Compared to the Coin Game or simple matrix games, the grid-world setting presents a variety of new challenges such as partial observability, multi-step agent interactions, temporally-extended actions, and longer time horizons. Unlike the “in the Matrix” games from Melting Pot, STORM features stochasticity, increasing the difficulty [15].
3.1.2 Existing Environments
We have also provided JAX-based implementations of several existing environments.
Hanabi [3] is a fully-cooperative partially-observable multiplayer card game, where players can observe others’ cards but not their own. It is a common benchmark for zero-shot coordination, theory of mind, and ad-hoc teamplay research [23, 24, 9, 39].
Overcooked is commonly used for assessing fully-cooperative and fully-observable Human-AI task performance. Our implementation mimics the original from Overcooked-AI [10]. For a discussion on this environment’s limitations see [33].
MABrax is a derivative of Multi-Agent MuJoCo [49], an extension of the MuJoCo Gym environment [62] that is commonly used for benchmarking continuous multi-agent robotic control.
Multi-Agent Particle Environment (MPE) tasks feature a 2D world with simple physics where particle agents can move, communicate, and interact with fixed landmarks [36].
Coin Game is a two-player grid-world environment which emulates social dilemmas such as the iterated prisoner’s dilemma [57]. While this is a common benchmark for the general-sum setting, previous work [29] has illustrated issues which STORM corrects.
Switch Riddle [16] is a simple cooperative communication task included as a debugging tool.
3.2 Algorithms
In this section, we present our re-implementation of five well-known MARL baseline algorithms using JAX.
All of our training pipelines are fully compatible with JAX’s jit and vmap functions, resulting in significant acceleration of the training processes, as outlined in Figure 3. It also enables parallelisation of training across many seeds and hyperparameters on a single GPU. We follow CleanRL’s philosophy of providing clear, single-file implementations [26] and provide a brief overview of the implemented baselines in the Appendix.
PPO
Q-learning
4 Evaluation Recommendations
Previous work [21] has found significant differences in the evaluation protocols between MARL research works. We identify four main research areas that would benefit from our library: cooperative centralised training with decentralised execution (CTDE) [16], zero-shot coordination [23], general-sum games, and cooperative continuous action methods.
| Setting | Recommended Environments |
| CTDE | SMAX (all scenarios), Hanabi (2-5 players), Overcooked |
| Cooperative Continuous Actions | MABrax |
To compute aggregate performance statistics, we follow the recommendations of [2] and evaluate the inter-quartile mean across the different classes of environments. To do this, we recommend normalising performance of the algorithms on the relevant classes of environment (for example via looking at the maximum and minimum performance across algorithms as discussed in [21]), and computing a mean per seed. Then compute the inter-quartile mean across aggregated statistics in each environment. We provide code for performing this calculation. This allows environment classes to be compared fairly, without over-weighting those with more individual scenarios.
5 Results
To demonstrate the utility of our library, we evaluate PPO against Q-Learning algorithms on a range of cooperative environments. We discover that not only does it have improved performance, but also that it is more practical to use for end-to-end GPU training.
We then evaluate the speed of our library. We compare algorithm and environment runtimes with similar CPU-based environments. We find that when training PPO, JaxMARL is 31x quicker on SMAX when compared to training in SMAC and 14x quicker on MPE for a single run, and 12,500x for vectorized training runs. Finally, we verify the correctness of our implementations by performing thorough comparisons with prior work.
5.1 Multi-Environment Comparison
We provide a preliminary comparison of our PPO and Q-Learning baselines in Figure 4. The IQM and mean were aggregated across 9 SMAX tasks, excluding the two maps with more than 10 units, all 5 Overcooked maps, and the 2 cooperative scenarios of MPE, running 10 seeds per task. We did not evaluate on Hanabi or all SMAX tasks because of the large memory overhead of storing the replay buffer for the Q-Learning methods on the GPU. We normalize the scores of each run on each task against the highest score obtained by any algorithm in that task, and then average the scores in each environment to avoid bias towards SMAX (which contributes more tasks).111For PPO, we used MAPPO as the baseline algorithm, except for Overcooked where we used IPPO. For Q-Learning, we used VDN, except for SMAX where we use QMIX as it performs slightly better.
| Environment | PPO | Q-Learning |
| SMAX | 10 | 60 |
| Overcooked | 1 | 10 |
| MPE | 1.5 | 2 |
| Hanabi | 450 | - |
5.2 Speed Benchmarking
We compare the performance of our environments in steps per second when using random actions to the original environments in Table 3, with details of this test provided in the Appendix.
| Environment | Original, 1 Env | Jax, 1 Env | Jax, 100 Envs | Jax, 10k Envs |
| MPE Simple Spread | ||||
| Switch Riddle | ||||
| Hanabi | ||||
| Overcooked | ||||
| MABrax Ant 4x2 | ||||
| Starcraft 2s3z | ||||
| Starcraft 27m vs 30m | ||||
| STORM | – | |||
| Coin Game |
We next compare the speed of our training pipeline to that of PyMARL. As shown in Figure 5(a), a single Q-Learning training run for MPE’s simple spread task takes 130 seconds with JaxMARL while PyMARL requires over an hour. Furthermore, using JAX we can parallelise over the entire training process within a single hardware accelerator. For QMIX on MPE, this allows us to complete 1024 individual training runs in 198.4 seconds, compared to 1 hour and 10 minutes for a single training run with PyMARL, a speed up of 21,500x per agent. This analysis is repeated for IPPO in the Appendix and we find a speedup of 12,500x. Figure Figure 5(c) demonstrates the speedup gained from using SMAX with JaxMARL’s IPPO implementation compared to training on SMAC with PyMARL. Across a varying number of environment rollout threads, JaxMARL gives a speedup of up to 31x.
5.3 Algorithm and Environment Correctness
In this section, we compare our environment and algorithm implementations with prior work and demonstrate equivalence where applicable.
Overcooked
The transition dynamics of our Overcooked implementation match those of the Overcooked-AI implementation. We demonstrate this by training an IPPO policy on our implementation and evaluating the policy on both our implementation and the original at regular intervals. The performance is similar across the implementations. Results can be found in the Appendix.
SMAX
SMAX and SMAC are different environments, as they have different opponent policies and dynamics. However, we demonstrate some similarity between them by comparing our IPPO and MAPPO implementations against MAPPO results on SMAC, using the implementation from [58]. We show this figure, along with a more in-depth description of their differences, in the appendix.
We additionally present aggregate and detailed performance across SMAX in the Appendix. The PPO-based IPPO and MAPPO perform better than the Q-Learning methods, with the centralised information provided in the state helping MAPPO significantly outperform IPPO.
MPE
Our MPE environment corresponds exactly to the PettingZoo implementation. We validate this for each environment using a uniform-random policy on rollouts, ensuring all observations and rewards are within a tolerance of at each transition. We additionally compare the results of our Q-Learning and PPO implementations with existing libraries, the results of which, along with the performance of IQL on the remaining MPE environments, can be found in the Appendix. We compare Q-Learning and PPO on Simple Spread in Figure 6(a).
MABrax
As Brax differs subtly from MuJoCo, MABrax does not correspond to MAMuJoCo but the learning dynamics are qualitatively similar. Results therefore are not directly comparable across the two environments.
We report mean training return across 10 seeds for IPPO on halfcheetah_6x1 in Figure 6(b). We additionally report the training curves for IPPO on ant_4x2, hopper_3x1, walker2d_2x3 and humanoid_9|8 in the Appendix.
Hanabi
Our implementation matches the Hanabi Learning Environment. To verify the environment’s accuracy, we obtained 10,000 action trajectories from the original C++ repository using their pretrained models. We confirmed that processing these action trajectories with JaxMARL produces the same states and returns as the C++ repository. In addition, we transferred the models trained with Pytorch/C++ to Jax and verified they obtain similar scores in JaxMARL. Finally, as shown in Figure 6(c), our IPPO and MAPPO models attain scores and respectively, which are better than (for IPPO) or similar to (for MAPPO) the scores attained in [67].
6 Related Work
MARL Libraries and Algorithms
Several open-source libraries exist for both MARL algorithms and environments. The popular library PyMARL [53] provides PyTorch implementations of QMIX, VDN and IQL and integrates easily with SMAC. E-PyMARL [48] extends this by adding the actor-critic algorithms MADDPG [36], MAA2C [44], IA2C [44], and MAPPO, and supports SMAC, Gym [8], Robot Warehouse [11], Level-Based Foraging [11], and MPE environments. Recently released MARLLib [25] is instead based on the open-source RL library RLLib [35] and combines a wide range of competitive, cooperative and mixed environments with a broad set of baseline algorithms. Meanwhile, MALib [70] focuses on population-based MARL across a wide range of environments. However, none of these frameworks feature hardware-accelerated environments and thus lack the associated performance benefits.
Hardware-Accelerated and JAX-Based RL
There has also been a recent proliferation of hardware-accelerated and JAX-based RL environments. Isaac gym [42] provides a GPU-accelerated simulator for a range of robotics platforms and CuLE [12] is a CUDA reimplementation of the Atari Learning Environment [4]. Both of these environments are GPU-specific and cannot be extended to other hardware accelerators. Madrona [56] is an extensible game-engine written in C++ that allows for GPU acceleration and parallelisation across environments. However, it requires environment code to be written in C++, limiting its accessibility. VMAS [5] provides a vectorized 2D physics engine written in PyTorch and a set of challenging multi-robot scenarios, including those from the MPE environment. For RL environments implemented in JAX, Jumanji [6] features mostly single-agent environments with a strong focus on combinatorial problems. The authors also provide an actor-critic baseline in addition to random actions. PGX [31] includes several board-game environments written in JAX. Gymnax [32] provides JAX implementations of the BSuite [47], classic continuous control, MinAtar [66] and other assorted environments. Gymnax’s sister-library, gymnax-baselines, provides PPO and ES baselines. Further extensions to Gymnax [40] also include POPGym environments [45]. Brax [19] reimplements the MuJoCo simulator in JAX and also provides a PPO implementation as a baseline. Jax-LOB [20] implements a vectorized limit order book as an RL environment that runs on the accelerator. Perhaps the most similar to our work is Mava [50], which provides a MAPPO baseline, as well as integration with the Robot Warehouse environment. None of these libraries combine a range of JAX-based MARL environments with both value-based and actor-critic baselines.
7 Conclusion
Hardware acceleration offers important opportunities for MARL research by lowering computational barriers, increasing the speed at which ideas can be iterated, and allowing for more thorough evaluation. We present JaxMARL, an open-source library of popular MARL environments and baseline algorithms implemented in JAX. We combine ease of use with hardware accelerator enabled efficiency to give significant speed-ups compared to traditional CPU-based implementations. Furthermore, by bringing together a wide range of MARL environments under one codebase, we have the potential to help alleviate issues with MARL’s evaluation standards. We hope that JaxMARL will help advance MARL by enabling researchers to conduct research with thorough, fast, and effective evaluations.
Limitations and Future Work.
While our work provides significant advancements, several limitations remain. First, we observe that the speedups are less pronounced for off-policy, value-based methods. Additionally, there are inherent challenges with end-to-end JAX implementations, such as the difficulty in efficiently handling environments with a variable number of agents or those with massive observation sizes. Furthermore, our MARL environments largely re-implement or draw inspiration from existing environment suites, meaning they do not yet push the boundaries of MARL capabilities. Developing novel MARL environments that push the boundaries of current capabilities could provide new and challenging benchmarks for the community.
8 Acknowledgements
This work received funding from the EPSRC Programme Grant “From Sensing to Collaboration” (EP/V000748/1). MG was partially founded by the FPI-UPC Santander Scholarship FPI-UPC_93. JF is partially funded by the UKI grant EP/Y028481/1 (originally selected for funding by the ERC). JF is also supported by the JPMC Research Award and the Amazon Research Award.
Appendix
We structure the appendix as follows. Appendix A provides further background on SMAC, Appendix B describes the environments included within JaxMARL, and Appendix C sets out JaxMARL’s API. Appendix D explores JaxMARL’s Value-Based MARL methods, including relevant implementation details. Appendix E reports our speed comparisons while Appendix F details training and correctness results not included in the main text. Our hyperparameter values are listed in Appendix G.
Appendix A Further Background on SMAC
StarCraft is a popular environment for testing RL algorithms. It typically features a centralised controller issuing commands to balance micromanagement, the low-level control of individual units, and macromanagement, the high level plans for economy and resource management.
SMAC [53], instead, focuses on decentralised unit micromanagement across a range of scenarios divided into three broad categories: symmetric, where each side has the same units, asymmetric, where the enemy team has more units, and micro-trick, which are scenarios designed specifically to feature a particular StarCraft micromanagement strategy. SMACv2 [15] demonstrates that open-loop policies can be effective on SMAC and adds additional randomly generated scenarios to rectify SMAC’s lack of stochasticity. However, both of these environments rely on running the full game of StarCraft II, which severely increases their CPU and memory requirements. SMAClite [43] attempts to alleviate this computational burden by recreating the SMAC environment primarily in NumPy, with some core components written in C++. While this is much more lightweight than SMAC, it cannot be run on a GPU and therefore cannot be parallelised effectively with typical academic hardware, which commonly has very few CPU cores compared to industry clusters.
Appendix B Further Details on Environments
B.1 SMAX